OpenStudy (anonymous):

An airplane flew 3 hours with a 60 mph head wind. The return trip with a tail wind of the same speed took 2 hours. Find the speed of the plane in still air.

OpenStudy (anonymous):

velocity (in MPH) * time (in hours) = distance (in miles) The distance is the same, it's EQUAL. Hint hint! The plane also is traveling at the same airspeed both ways Therefore: \[(velocity)(time)=(v-60)(3)=(v+60)(2)\]

OpenStudy (anonymous):

Thank You =]

OpenStudy (anonymous):

\[3 (v-60) = 2 (v+60)\] Expand out terms on both sides: \[3 v-180 = 2 v+120\] Add (180-2 v) to both sides: \[v = 300\] That's 300MPH airspeed, the distance is 720 miles I believe. Check me :-)